# Mining Limited Data Sufficiently: A BERT-inspired Approach for CSI Time Series Application in Wireless Communication and Sensing

### v1

@@ IV-C3 CSI Classification Task
The robust and powerful capabilities of our CSI-BERT2 model in CSI classification tasks are demonstrated in Table III, where it achieves the best performance across all scenarios. Moreover, with the aid of time embedding, our CSI-BERT2 can be employed in scenarios where the training set has samples with different sampling rates, or the testing set has a different sampling rate from the training set. This is a practical feature, especially in federated learning scenarios, where we cannot guarantee that all clients have the same sampling rate. In Table V, we compare the performance of different models where the training and testing sets both have 100Hz and 50Hz CSI, as well as where the training set only has 100Hz CSI but the testing set has 50Hz CSI. For fairness, we provide the CSI processed by linear interpolation for the other non-CSI-BERT models. It can be seen that other models are significantly influenced by the different sampling rates, with some even failing in particular scenarios. In contrast, CSI-BERT2 is almost unaffected by any impact, thanks to the help of its time embedding.
What’s more, resampling methods could also be a potential way to address such scenarios. However, on one hand, the resampling operation may introduce a relatively high time delay. On the other hand, these resampling methods were not designed specifically for CSI, which means they may not be able to capture the inherent relationships within the CSI as traditional interpolation methods. The resampling error could also potentially harm the model’s performance.
@@
Data Training Set: 100Hz+50Hz; Testing Set: 100Hz+50Hz Model MLP CNN RNN [47] LSTM [45] Chen et al. [48] WiGRUNT [46] CSI-BERT1 [24] CSI-BERT2 Gesture Recognition 16.51% 14.54% 15.13% 17.37% 16.59% 74.47% 64.61% 97.04% People Identification 13.47% 15.52% 13.41% 13.47% 13.72% 81.25% 70.83% 99.54% Action Recognition 70.97% 67.67% 70.02% 59.63% 75.18% 66.78% 78.18% 88.35% Fall Detection 81.10% 75.80% 83.92% 84.83% 85.98% 80.72% 92.98% 93.64% People Number Estimation 84.03% 46.82% 81.25% 82.00% 80.09% 76.13% 86.93% 92.54% Data Training Set: 100Hz; Testing Set: 50Hz Model MLP CNN RNN [47] LSTM [45] Chen et al. [48] WiGRUNT [46] CSI-BERT1 [24] CSI-BERT2 Gesture Recognition 69.79% 20.38% 36.25% 27.15% 74.89% 71.37% 79.96% 97.81% People Identification 87.29% 11.85% 82.29% 22.32% 87.22% 85.24% 94.44% 99.38% Action Recognition 68.97% 51.96% 68.07% 60.26% 76.56% 73.21% 84.56% 88.53% Fall Detection 80.83% 76.13% 80.17% 84.38% 78.61% 77.08% 94.19% 94.32% People Number Estimation 82.22% 77.03% 84.43% 42.22% 68.89% 71.11% 89.10% 94.77%
@@ IV-D Ablation Study
IV-D1 Effect of Pre-training
To illustrate the help of our pre-training mechanism for downstream tasks, we compare the model’s performance with and without pre-training, as shown in Fig. 7 and Table VI. In the CSI classification task, the improvement from pre-training is modest but not particularly significant. However, in the CSI prediction task, pre-training proves to be highly beneficial, likely because the prediction task is more closely related to the mask recovery task.
@@ IV-D1 Effect of Pre-training
w/ Pre-training w/o Pre-training Dataset MSE SMAPE MAPE MSE SMAPE MAPE WiGesture [24] 3.2942 0.1583 0.1349 5.3054 0.1962 0.1657 WiFall [26] 4.8598 0.1471 0.1347 5.0957 0.1595 0.1413 Wicount 5.4301 0.1726 0.1590 6.6868 0.2019 0.1659
IV-D2 Effect of Change on Original BERT Model
To demonstrate the effectiveness of the modifications we made to BERT [20], we first compare the recovery performance of the original BERT (where we simply use a linear layer as the embedding and output layer to support the CSI format) with our modified version, as shown in Fig. 8. We observe that the amplitude spectrum of the original BERT appears smooth across all subcarriers, indicating that it tends to map all tokens within each subcarrier to similar value. Although BERT can also achieve a relatively low MSE of 5.26 in WiGesture dataset, it fails to capture any valuable information.
@@ IV-D2 Effect of Change on Original BERT Model
Next, we compare the attention patterns of the proposed CSI-BERT models with the original BERT. As shown in Fig. 9, we illustrate the mean of attention patterns across all attention heads. The attention patterns of CSI-BERT1 and CSI-BERT2 consist of multiple diagonals from left to right, which is interpretable. This is because in a short time interval, CSI always changes in a particular regular pattern. Additionally, it is reasonable that their attention patterns are approximately symmetric, which corresponds to the mechanism of bi-directional attention in BERT. However, we cannot find any regular patterns in the attention patterns of the original BERT.
(a) BERT [20]
(sample set 1)
(b) CSI-BERT1 [24]
(sample set 1)
(c) CSI-BERT2
(sample set 1)
(d) BERT [20]
(sample set 2)
(e) CSI-BERT1 [24]
(sample set 2)
(f) CSI-BERT2
(sample set 2)
@@
Furthermore, we use a Gradient-weighted Class Activation Mapping (Grad-CAM) [60] method to provide explanations for the CSI recovery task of CSI-BERT and BERT. We manually delete a given package and use the models to recover them. After that, we calculate the Grad-CAM based on the gradient of the original input. We then combine it with the original input as shown in Fig. 10, which represents what the model mainly focuses on when recovering the deleted package. It can be seen that our CSI-BERT1 and CSI-BERT2 models mainly focus on packages nearby the lost package, which is consistent with our intuition. However, the original BERT only focuses on the lost package, which is actually meaningless because the lost package is replaced by [MASK] and can hardly provide useful information, except that the package at this position is lost.
(a) BERT [20]
($20^{th}$ package)
(b) CSI-BERT1 [24]
($20^{th}$ package)
(c) CSI-BERT2
($20^{th}$ package)
(d) BERT [20]
($80^{th}$ package)
(e) CSI-BERT1 [24]
($80^{th}$ package)
(f) CSI-BERT2
($80^{th}$ package)
Figure 10: Grad-CAM: The two rows of heatmaps illustrate the Grad-CAM when we manually delete a package at position 20 and position 80. The blank positions represent the actual lost CSI during transmission. We calculate Grad-CAM solely based on the loss function at the positions where we manually delete the packages.
@@ V Conclusion
In this paper, we introduce CSI-BERT2, a multifunctional model designed for CSI time series, encompassing CSI prediction, and classification. Through the proposed pre-training and fine-tuning methods, CSI-BERT2 effectively extracts valuable insights from limited CSI data. We have developed a novel time embedding mechanism to enhance the suitability of Transformer-based models for time series analysis. Additionally, we employed an ARL model to improve the model’s ability to capture relationships between different subcarriers. We also propose the MPM task, which fine-tunes BERT-based models specifically for prediction tasks. Furthermore, we have optimized the training process of CSI-BERT1, resulting in a significant enhancement in the model’s performance.
Through extensive experimentation, we demonstrate the effectiveness of CSI-BERT2. In the CSI prediction task, it significantly outperforms traditional models while maintaining relatively fast computation speeds. In the CSI classification task, it achieves state-of-the-art performance and can effectively handle scenarios where the training and testing sets contain samples with different sampling rates.
The CSI-BERT2 model can be further applied in other CSI-related scenarios and time-series tasks. Moreover, we aim to explore how to leverage multi-source heterogeneous CSI data to train our CSI-BERT simultaneously. By doing so, we hope to replicate the generalization capabilities of Large Language Models (LLMs) and Large Vision Models (LVMs) in the context of wireless sensing.

### v2

@@ IV-C3 CSI Classification Task
The robust and powerful capabilities of our CSI-BERT2 model in CSI classification tasks are demonstrated in Table III, where it achieves the best performance across all scenarios. Notably, the original CSI-BERT fails to outperform models such as Chen et al. [51] and WiGrumt [49] in many scenarios. However, our CSI-BERT2 shows substantial improvement due to the enhanced model structure, thanks to the ARL capturing the relationships and importance among subcarriers, and the MLP-based time embedding layer that encodes time information more effectively. Moreover, aided by the time embedding mechanism, CSI-BERT2 can be effectively applied in scenarios where the training data includes samples with varying sampling rates, or where the sampling rate of the testing set differs from that of the training set. This is a practical feature, especially in federated learning scenarios, where we cannot guarantee that all clients have the same sampling rate. In Table V, we compare the performance of various models under three conditions: (1) both training and testing sets contain 100 Hz CSI, (2) both contain 50 Hz CSI, and (3) the training set contains only 100 Hz CSI while the testing set contains 50 Hz CSI. For fairness, we provide the CSI processed by linear interpolation for the other non-CSI-BERT models. It can be seen that other models are significantly influenced by the different sampling rates, with some even failing to function properly in particular scenarios. In contrast, CSI-BERT2 is almost unaffected by any impact, thanks to the help of time embedding.
What’s more, resampling methods could also be a potential way to address such scenarios. However, on one hand, the resampling operation may introduce a relatively large delay. On the other hand, these resampling methods were not designed specifically for CSI, which means they may not be able to capture the inherent relationships within the CSI as traditional interpolation methods. Therefore, the resampling error could also potentially harm the model’s performance.
@@
Data Training Set: 100Hz+50Hz; Testing Set: 100Hz+50Hz Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Gesture Recognition 16.51% 14.54% 15.13% 17.37% 16.59% 74.47% 64.61% 97.04% People Identification 13.47% 15.52% 13.41% 13.47% 13.72% 81.25% 70.83% 99.54% Action Recognition 70.97% 67.67% 70.02% 59.63% 75.18% 66.78% 78.18% 88.35% Fall Detection 81.10% 75.80% 83.92% 84.83% 85.98% 80.72% 92.98% 93.64% People Number Estimation 84.03% 46.82% 81.25% 82.00% 80.09% 76.13% 86.93% 92.54% Data Training Set: 100Hz; Testing Set: 50Hz Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Gesture Recognition 69.79% 20.38% 36.25% 27.15% 74.89% 71.37% 79.96% 97.81% People Identification 87.29% 11.85% 82.29% 22.32% 87.22% 85.24% 94.44% 99.38% Action Recognition 68.97% 51.96% 68.07% 60.26% 76.56% 73.21% 84.56% 88.53% Fall Detection 80.83% 76.13% 80.17% 84.38% 78.61% 77.08% 94.19% 94.32% People Number Estimation 82.22% 77.03% 84.43% 42.22% 68.89% 71.11% 89.10% 94.77%
@@ IV-D Ablation Study
IV-D1 Effect of Unsupervised Training
To illustrate the benefits of our unsupervised training mechanism for downstream tasks, we compare the model’s performance with and without unsupervised training, as shown in Fig. 6 and Table VI. While unsupervised training contributes to performance gains in the CSI classification task, the improvement remains relatively modest. However, in the CSI prediction task, unsupervised training proves to be highly beneficial, this is because the prediction task is more closely related to the mask recovery task.
@@ IV-D1 Effect of Unsupervised Training
Dataset w/ Unsupervised Training w/o Unsupervised Training
MSE SMAPE MAPE MSE SMAPE MAPE
CommPre 1.5689 ×$\times 10^{-8}$ 0.2959 0.2319 2.0080 ×$\times 10^{-8}$ 0.3198 0.2477
WiGesture [7] 3.2942 0.1583 0.1349 5.3054 0.1962 0.1657
WiFall [37] 4.8598 0.1471 0.1347 5.0957 0.1595 0.1413
WiCount 5.4301 0.1726 0.1590 6.6868 0.2019 0.1659
IV-D2 Effect of Changes on Original BERT Model
To demonstrate the effectiveness of the modifications we made on BERT [18], we first compare the recovery performance of the original BERT (where we simply use a linear layer as the embedding and output layer to support the CSI format) with our modified version, as shown in Fig. 7. We observe that the amplitude spectrum of the original BERT appears smooth across all subcarriers, indicating that it tends to map all tokens within each subcarrier to similar value. Although BERT can also achieve a relatively low MSE of 5.26 in WiGesture dataset, it fails to capture any valuable information.
@@ IV-D2 Effect of Changes on Original BERT Model
Next, we compare the attention patterns of the proposed CSI-BERT models with the original BERT. As shown in Fig. 8, we illustrate the mean of attention patterns across all attention heads. The attention patterns of CSI-BERT and CSI-BERT2 consist of multiple diagonals from left to right, which is interpretable. This is because in a short time interval, CSI always changes in a particular regular pattern. Additionally, it is reasonable that their attention patterns are approximately symmetric, which corresponds to the mechanism of bi-directional attention in BERT. However, we cannot find any regular patterns in the attention patterns of the original BERT.
Furthermore, we use a gradient-weighted class activation mapping (Grad-CAM) [59] method to provide explanations for the CSI recovery task of CSI-BERT and BERT. We manually delete a given packet and use the models to recover them. After that, we calculate the Grad-CAM based on the gradient of the original input. We then combine it with the original input, as illustrated in Fig. 9, to visualize the regions the model primarily attends to during the recovery of the missing packet. It can be seen that our CSI-BERT and CSI-BERT2 models mainly focus on packets nearby the lost packet, which is consistent with our intuition. However, the original BERT only focuses on the lost packet, which is actually meaningless because the lost packet is replaced by [MASK] and can hardly provide useful information, except that the packet at this position is lost.
Finally, by comparing the attention maps and Grad-CAM between CSI-BERT and CSI-BERT2, we can explain why CSI-BERT2 exhibits higher generalization. In Fig. 9, we observe that CSI-BERT2 has a more centralized activation map, suggesting it utilizes less information for training while achieving better performance than CSI-BERT. This allows CSI-BERT2 to learn a simpler attention map, as shown in Fig. 8, enabling it to focus on more robust features rather than overfitting to specific relationships in the training set. This improved generalization may be attributed to the ARL and time embedding layers, which help capture more information in the dimensions of subcarrier and time (packet).
(a) BERT [18] (sample set 1)
(b) CSI-BERT [7] (sample set 1)
(c) CSI-BERT2 (sample set 1)
(d) BERT [18] (sample set 2)
(e) CSI-BERT [7]  (sample set 2)
(f) CSI-BERT2 (sample set 2)
@@
(a) BERT [18] (202020-th packet)
(b) CSI-BERT [7] (202020-th packet)
(c) CSI-BERT2 (202020-th packet)
(d) BERT [18] (808080-th packet)
(e) CSI-BERT [7] (808080-th packet)
(f) CSI-BERT2 (808080-th packet)
Figure 9: Grad-CAM: The two rows of heatmaps illustrate the Grad-CAM when we manually delete a packet at position 20 and position 80. The blank positions represent the actual lost CSI during transmission. We calculate Grad-CAM solely based on the loss function at the positions where we manually delete the packets.
@@ V Conclusion
In this paper, we have proposed CSI-BERT2, a multifunctional model designed for CSI time series tasks, including both CSI prediction, and classification. Leveraging the proposed two-stage training framework, CSI-BERT2 is capable of extracting meaningful representations from limited CSI data. We have developed a novel time embedding mechanism to enhance the suitability of Transformer-based models for time series analysis. Additionally, we have employed an ARL model to improve the model’s ability to capture relationships among subcarriers. Moreover, we have proposed an MPM that fine-tunes the BERT architecture specifically for CSI prediction. Furthermore, we have optimized the original training process of CSI-BERT, leading to a significant enhancement in the model’s performance. These novel designs successfully address the low performance of CSI-BERT in CSI classification tasks and expand its application scenario to include CSI prediction, all within a unified framework. Extensive experiments have validated the effectiveness of CSI-BERT2. In the CSI prediction task, it significantly outperforms traditional models while maintaining relatively fast computation speeds. In the CSI classification task, it achieves SOTA performance, even under challenging conditions where training and testing sets differ in sampling rates
The CSI-BERT2 model can be further applied in other CSI-related scenarios and time-series tasks. Moreover, we aim to explore how to leverage multi-source heterogeneous CSI data to train our CSI-BERT2 simultaneously. By doing so, we hope to replicate the generalization capabilities of large language models (LLMs) and large vision models (LVMs) in the context of wireless sensing.